In a competitive market the demand curve is given by Q = 1600 –4P and supply by Q = 4P. What is the total surplus (TS) in the market equilibrium?

a) TS = 80,000

b) TS = 1,600

c) TS = 0

d) TS = 160,000

e) none of the above.
Answer #1

Qd = 1600- 4P
When P=0 , Q=1600

When Q=0 , P=400

Qs = 4P

When Q=0 , P=0

At equilibrium : Qd = Qs

1600-4P = 4P

1600= 8P

P= $200 (Equilibrium price)

Q = 4P= 4(200)= 800 units (Equilibrium quantity)

Consumer surplus = (1/2)(400-200)(800)= $80000

Producer surplus = (1/2)(200-0)(800)= $80000

Total surplus = CS+PS = $(80000+80000)= $160,000.

Hence, option(D) is correct.